# Hypercube Policy Regularization Framework for Offline Reinforcement Learning

### v1

@@ Abstract
Offline reinforcement learning has received extensive attention from scholars because it avoids the interaction between the agent and the environment by learning a policy through a static dataset. However, general reinforcement learning methods cannot get satisfactory results in offline reinforcement learning due to the out-of-distribution state actions that the dataset cannot cover during training. To solve this problem, the policy regularization method that tries to directly clone policies used in static datasets has received numerous studies due to its simplicity and effectiveness. However, policy constraint methods make the agent choose the corresponding actions in the static dataset. This type of constraint is usually over-conservative, which results in suboptimal policies, especially in low-quality static datasets. In this paper, a hypercube policy regularization framework is proposed, this method alleviates the constraints of policy constraint methods by allowing the agent to explore the actions corresponding to similar states in the static dataset, which increases the effectiveness of algorithms in low-quality datasets. It was also theoretically demonstrated that the hypercube policy regularization framework can effectively improve the performance of original algorithms. In addition, the hypercube policy regularization framework is combined with TD3-BC and Diffusion-QL for experiments on D4RL datasets which are called TD3-BC-C and Diffusion-QL-C. The experimental results of the score demonstrate that TD3-BC-C and Diffusion-QL-C perform better than state-of-the-art algorithms like IQL, CQL, TD3-BC and Diffusion-QL in most D4RL environments in approximate time.
@@
reinforcement learning , Offline reinforcement learning , policy regularization
journal: Neural Networks
[…]
[university]organization=Sun Yat-Sen University,addressline=135 Xingang West Road, Guangzhou, city=Guangzhou, postcode=510275, state=Guandong, country=China
@@ 1 Introduction
The objective of Reinforcement Learning(RL) is to learn an optimal policy through the interaction between the agent and the environment. Given the similarity between the manner of RL to generate diverse policies and cognitive abilities, it has received extensive attention and research. At present, there have been some notable achievements in the field, such as games[23, 27], medicine[32],sport[5] and navigator[12].
@@
However, a potential issue with general RL is that the agent must interact with the environment during training, which may cause harm to the environment or the agent, leading to economic losses[3]. One potential solution to this problem is offline Reinforcement learning (offline RL)[19]. In contrast to RL, in offline RL, the agent no longer learns a policy through interaction with the environment. Instead, the agent tries to learn an optimal policy from a static dataset. Since the training of offline RL only requires a static dataset, it does not have the problem of the environment or agent being damaged during training. Nevertheless, offline RL only learns from a limited static dataset, which is typically unable to cover the full spectrum of potential state-action combinations. These state-action pairs that are not included in the static dataset are called out-of-distribution (OOD) state actions[8]. Function estimators employed during offline RL training are prone to inaccuracy when estimating the state-action value (Q-value) or state value (V-value) of an OOD state action. This erroneous overestimation of some Q-values or V-values is likely to result in the agent selecting some suboptimal state-actions during training, This preference for suboptimal state-actions ultimately leads to the inability to obtain a satisfactory policy. To address this issue, offline RL has developed a set of techniques that are distinct from those employed in general RL.
@@
Current offline RL methods can be mainly classified into policy regularization[21, 31], Q-value regularization[1, 22]. importance sampling[11, 34], uncertainty estimation[30, 4] and model-based methods[13, 14, 20, 33]. Among the aforementioned methods, Q-value regularization methods constrain the Q-value of OOD state-actions, thereby obtaining a policy through the exploration of the constrained state action. Policy regularization methods attempt to clone the policy used for static datasets. Current algorithms utilizing policy regularization methods can achieve good results by using it in conjunction with advanced policy clone models such as diffusion models to clone the behavior of policy distribution[28]. Current Q-value regularization methods use ensemble[4], count-based[14], and other methods to more accurately constrain the Q-value of OOD state actions, thereby facilitating enhanced exploration.
The motivation behind this article is as follows:
(1) Policy regularization: One advantage of policy regularization methods is that they have the potential to get a good policy in a shorter time than other methods[7]. However, when faced with low-quality dataset, policy regularization methods will also endeavor to replicate the bad actions within it, resulting in a suboptimal policy[18]. A potential solution to this problem may be found by undertaking an examination of some relevant actions.
(2)Q-value regularization: Q-value regularization methods can perform well in low-quality datasets by exploring the whole space. However, if the agent is allowed to explore the whole space, it is necessary to constrain the Q-value of OOD state-actions, which will result in a significant increase in the number of hyperparameters and operation time. A potential solution to this problem may be found in the reduction of the exploration range of the agent.
@@
(3)Improvement idea: Due to the aforementioned issues, the combination of policy regularization and Q-value regularization methods may yield superior results. Therefore, this article seeks to permit the agent to explore some promising actions and replicate the overall state actions that have been explored. It can be noted that for any given state, there may be multiple similar states within the static dataset. Furthermore, in datasets, especially low-quality datasets, such similar states may have better actions due to the randomness of the sampling actions. Therefore, an attempt is made to set the exploration scope of the agent in a given state limited to actions that correspond to adjacent states.
@@
In light of the above ideas, motivated by Latin hypercube sampling[24], which transforms the space into hypercubes, the hypercube policy regularization framework is proposed. In this framework, the state space is divided into hypercubes and the agent is permitted to explore the actions corresponding to the states in the static dataset within the same hypercube. Compared with current policy regularization methods, which are ineffective in low-quality datasets, and the current Q-constraint methods, which require a long training time[17]. The hypercube policy regularization framework enables agents to explore actions to a certain extent while maintaining sufficient constraints on out-of-distribution state actions, thereby improving the performance of algorithms while maintaining the same training time as policy regularization methods. Another advantage of this framework is its versatility, which allows it to be applied to a variety of policy regularization algorithms such as TD3-BC[7] and Diffusion-QL[28]. The combined algorithm, TD3-BC-C and Diffusion-QL-C exhibited remarkable performance in experiments conducted on the D4RL dataset[6]. TD3-BC-C and Diffusion-QL-C can achieve better results in a similar time compared to those of state-of-the-art(SOTA) algorithms such as Diffusion-QL and IQL[15].
@@
The remainder of this paper is organized as follows. Firstly, the related work and the required basic knowledge of this article are introduced. Secondly, a detailed explanation of the hypercube policy regularization framework is provided. Finally, experiments for the new algorithms are conducted on the D4RL dataset.
@@ 2 Related Work
This article is mainly related to policy regularization. In offline RL, The basic idea of the policy regularization method is to solve the OOD state action problem by penalizing the state action deviation from the static dataset with the distance between the trained action state and the corresponding state action within the static dataset. Among the policy constraint methods, BCQ[8] considers the extrapolation error and adds disturbances based on the use of VAE to generate actions to solve the extrapolation error. Therefore, BCQ can directly learn a good policy from a static dataset. BEAR[16] considers that previous algorithms cannot fit out-of-distribution(OOD) state-actions well. To deal with it, BEAR adds weighted behavior clone loss in the optimization process of the policy to solve the problem of OOD state actions by minimizing Maximum Mean Discrepancy(MMD)[10]. Compared with BCQ and BEAR, which require more operations, TD3-BC[7] achieves better results by adding only a few modifications to BC. TD3+BC adds behavior clone regularization terms to the value function and normalizes the state, using techniques similar to BEAR to achieve better results and less training time. BRAC[29] evaluates various methods of behavior clone regularization and proposes to use value penalty and policy regularization to improve algorithm efficiency. Compared with the above algorithms, IQL [15] additionally introduces the state value function to update the Q-value function and policy. Diffusion-QL[28] achieves a state-of-the-art(SOTA) level by using conditional diffusion models for behavior clones and policy regularization. These methods have achieved good results by modifying the behavior cloning model or adding regularization terms. However, as mentioned before, policy regularization methods rely on behavior clones of data within static datasets, which may result in suboptimal results when the quality of static datasets is poor.
@@
On the other hand, advanced algorithms constrain the Q-values of OOD state-actions and thus allow the agent to explore the entire state-action space. For instance, CQL[17], EDAC[2] and PBRL [4] can also have good results in low-quality datasets, but such algorithms usually require large computing time and computing space to find an excellent policy, and such algorithms cannot solve hard tasks well. Considering the characteristics of the distinguishing features of the two methods, local space exploration is integrated into the overall policy constraints. The goal is to enhance the performance of policy regularization methods in low-quality datasets while maintaining the capabilities of policy regularization methods.
3 Preliminaries
Offline RL: In general reinforcement learning, the environment is defined by a Markov decision process as $(S,A,r,P,{\rho_{0}},\gamma)$[25]. Here, $S$ is the state space, $A$ is the action space, $r:S\times A\to[R_{\min},{R_{\max}}]$ is the reward function, $P:S\times A\to S^{\prime}$ is the transition dynamics. By $P$, the agent translates to a new state after the current state, $\rho_{0}\in S$ is the initial state distribution and $\gamma\in(0,1]$ is the discount factor, which is used to adjust the impact of state-actions over time. The goal of RL is to learn a policy $\pi:S\to A$, by maximizing the expected return:
${E_{a_{i}\sim\pi}[\sum_{i=0}^{T-1}\gamma^{i}r(s_{i},a_{i})]}$
over a finite time horizon $T$. In offline RL, the environment is replaced by a static dataset defined as $D=\{(s,a,r,s^{\prime})\}$, here $s$ are the states, $a$ are the actions and $s^{\prime}$ are the next state after actions collected by a policy $\pi$ and the agent no longer interact with the environment to learn a policy but completely learns from the static dataset. The learning objectives and other parts are similar to general RL.
Policy constraint method: In offline RL, the policy constraint method represents a significant approach, This method obtains a policy by constraining the learned policy to align with specific characteristics of the behavior policy. The most fundamental policy constraint method directly obtains a policy through the least squares estimation, it can be expressed as follows:
$\pi=arg{\min}E_{({s_{i}},{a_{i}})\sim D}[\sum_{i=0}^{T-1}(\pi_{\theta}(s_{i})-a_{i})^{2}]$ (1)
Where $\pi_{\theta}$ is the training policy determined by parameter $\theta$, and $a_{i}$ is the action that follows this policy distribution[26]. The above equation directly implements behavior cloning by minimizing the difference between the policy-selected actions and the actions in the static dataset. Subsequent policy regularization methods use different methods to better clone the policy.
@@ 4 Proposed Method
In this section, the hypercube policy regularization framework is proposed and a theoretical analysis of the conditions under which the framework can ensure an enhancement in the performance of algorithms is offered. Finally, a new algorithm is created by combining the hypercube policy regularization framework with a baseline offline RL algorithm.
@@ 4.1 Hypercube policy regularization framework
The Q-function approximator is capable of accurately estimating the Q-value of state actions $(s,a)$ in static datasets. It is therefore proposed that the Q-function approximator can also be used to accurately estimate the Q-value of $(s^{\prime},a)$ with $|s-s^{\prime}|\leq\varepsilon$. Given these considerations, the hypercube policy regularization framework is proposed.
Figure 1: Hypercubed state space of dimension 3, choosing $\delta=2$.
The hypercube policy regularization framework is a method within the domain of offline RL that enables the agent to explore a range of actions within a hypercube, subsequently utilizing the best actions within this hypercube for policy clone. Compared with general policy regularization methods that only use the corresponding actions in the static dataset for policy clone during training, the hypercube policy regularization framework allows the agent to explore actions in the dataset that correspond to some similar states during the training process. Subsequently, the best action can be employed for policy clones, thereby enhancing the performance of the algorithm.
The hypercube partitioning can be obtained through the following two steps. The initial step is to divide the state space in the static dataset into a hypercube using an integer $\delta$. This is achieved by obtaining the coordinates of the $i$-th dimension of any state $s$ in the static dataset in hypercube space, as outlined below:
@@
Figure 1 shows a hypercubed state space of dimension 3, the $\delta=2$. Each circle represents a distinct state within the static dataset. The number of states within each hypercube $\geq 0$. During the training phase, the agent is permitted to explore the actions corresponding to the states in the same cube.
@@
In detail, in any state, the state-action value function $Q(s,a_{\max})$ of selecting the current optimal action within the same hypercube is compared with the state-action value function $Q(s,a)$ of selecting the action corresponding to that state itself, in order to select the better action and achieve the goal of enabling the agent to explore within a certain range. Finally, the hypercube policy regularization framework can be derived by integrating the constraint with a Markov decision process. Figure 2 illustrates a training flow chart that excludes the evaluation phase. At the beginning of training, the static dataset is hypercubed and it was used to constrain the current action $a$ to obtain a new action $a_{new}$ during training. At the end of a loop, the process is reversed to update the policy $\pi_{\theta}$, and this step is repeated to output a policy $\pi_{\theta}$.
4.2 Superiority analysis
If the Q-function approximator provides an entirely correct estimate of the Q-value, then it follows that utilizing the hypercube policy regularization framework can maintain or enhance the performance of the algorithm under any circumstances. However, in practical applications, the issue of imprecise estimation of the Q-value must be taken into account. When one state in a hypercube takes the optimal action corresponding to another state within the same hypercube, the Q-value corresponding to the state action $(s,a)$ may be overestimated. It is demonstrated that modifying the value of $\delta$ allows the hypercube constraint framework to enhance or sustain the performance of the algorithm regardless of the precision of the Q-value estimation.
Assumption 1.
@@ 4.2 Superiority analysis
$\left|{Q({s_{1}},{a_{1}})-Q({s_{2}},{a_{2}})}\right|\leq K\left|{({s_{1}}\oplus{a_{1}})-({s_{2}}\oplus{a_{2}})}\right|$ (5)
Where $\oplus$ is the vector concatenation operation, this assumption can be satisfied when the Q-function is approximated by neural networks[9]. Thus, we demonstrate that the hypercube policy regularization framework ensures that the algorithm’s performance does not deteriorate under Assumption 1.
Assumption 2.
Assuming that the Q-function approximator can accurately estimate the Q-value of state-actions within the static dataset.
Since the Q-function approximator estimates the Q-values of all state-actions through the data in the static dataset, It seems reasonable to posit that the Q-function approximation is capable of providing an accurate estimate of the Q-value of state-actions within the static dataset.
Theorem 1.
@@
$Q(s,{a^{\pi_{new}}})\geq Q(s,{a^{\pi_{old}}})$ (6)
Proof.
When only the $(s,a)$ exists in the hypercube, then $\pi_{new}=\pi_{old}$ and $Q(s,a^{\pi_{old}})=Q(s,a^{\pi_{new}})$.
When there is another state-action $(s^{\prime},a^{\prime})$ in the hypercube, according to Assumption 2, there is an accurate estimate of the Q-value relationship between $(s,a)$ and $(s^{\prime},a^{\prime})$. Therefore, $Q(s,a^{\prime})$ is discussed through $Q(s,a)$ and $Q(s^{\prime},a^{\prime})$. For this, three cases must be considered: $Q(s,a)<Q(s^{\prime},a^{\prime})$, $Q(s,a)=Q(s^{\prime},a^{\prime})$ and $Q(s,a)>Q(s^{\prime},a^{\prime})$.
When $Q(s,a)>Q(s^{\prime},a^{\prime})$, $Q(s,a)$ can be expressed as:
@@
$\begin{array}[]{c}Q(s,a^{\prime})\leq Q(s^{\prime},a^{\prime})+K|(s\oplus a^{\prime})-(s^{\prime}\oplus a)|\\ =Q(s^{\prime},a^{\prime})+K|s-s^{\prime}|\end{array}$ (8)
combing Equation(7) and inEquation(8), it can be known that when $K|s-s^{\prime}|<b_{1}$ , $Q(s,a^{\prime})\leq Q(s,a)$.
@@
This guarantees that when $\theta>c_{1}$, $Q(s,a)>Q(s^{\prime},a^{\prime})$, the new policy is always worse than the original policy, so $\pi_{new}=\pi_{old}$, $Q(s,a^{\pi_{new}})=Q(s,a^{\pi_{old}})$ in this situation.
When $Q(s,a)=Q(s^{\prime},a^{\prime})$, $\pi_{new}=\pi_{old}$ is selected to obtain a conservative policy. Therefore, in this case, $Q(s,a^{\pi_{new}})=Q(s,a^{\pi_{old}})$.
@@
$\begin{array}[]{c}Q(s,a^{\prime})\geq Q(s^{\prime},a^{\prime})-K|(s\oplus a^{\prime})-(s^{\prime}\oplus a^{\prime})|\\ =Q(s^{\prime},a^{\prime})-K|s-s^{\prime}|\end{array}$ (11)
Combining Equation(10) and inEquation(11), it can be known that when $K|s-s^{\prime}|<b_{2}$, $Q(s,a^{\prime})\geq Q(s,a)$. Similar to $Q(s^{\prime},a^{\prime})<Q(s,a)$, there exists an integer $c_{2}$, when $\theta>c_{2}$,
$K|s-s^{\prime}|\leq KS_{max}\leq b_{2}$ (12)
This demonstrates that when $\theta>c_{2}$, $Q(s^{\prime},a^{\prime})>Q(s,a)$. Therefore, regarded $a^{\prime}$ as $a_{new}$, the new policy can guarantee that $Q(s,a^{\pi_{new}})\geq Q(s,a^{\pi_{old}})$.
For the situation that there is only one extra $(s^{\prime},a^{\prime})$ in the hypercube and $\theta>max\{c_{1},c_{2}\}$. When $Q(s^{\prime},a^{\prime})\leq Q(s,a)$, the policy remains unchanged, $Q(s,a^{\pi_{new}})=Q(s,a^{\pi_{old}})$. When $Q(s^{\prime},a^{\prime})>Q(s,a^{\prime})$, $Q(s,a^{\pi_{new}})\geq Q(s,a^{\pi_{old}})$. Therefore, inEquation(6) holds in this case.
When there are multiple state-actions $(s_{i},a_{i}),i=1,2,...,n$ within the static dataset in each hypercube, refer to the above analysis of the existence of only one additional state-action in the hypercube, it can be seen that When the Q-value of some state-actions is smaller than $Q(s,a)$, there exist $b_{1}>0$, $\theta>c_{1}$ make:
@@
Therefore, when there are multiple state-actions in a hypercube, selecting $\theta>max\{c_{1},c_{2}\}$ can also make $Q(s,{a^{\pi_{new}}})\geq Q(s,{a^{\pi_{old}}})$.
This means that the hypercube policy regularization framework guarantees that the performance of the algorithm is not diminished, thereby enabling the framework to be applied in many scenarios.
In theory, selecting a large value for the parameter $\delta$ can ensure that the performance of the algorithms remains unaltered or improves. In practice, the Q-function approximator is capable of producing a relatively accurate estimate of the explored state actions within the static dataset, Consequently, the uncertainty associated with the Q-value obtained from exploring actions within the static dataset is likely to be relatively low. Therefore, to improve the effectiveness of the algorithms, smaller $\delta$ should be used to make the agent explore more actions so that the agent can maintain a balance between conservatism and aggressiveness and thus achieve the best results.

### v2

# Improving Offline Reinforcement Learning with Hypercube-Enhanced Policy Regularization

@@ Abstract
Offline reinforcement learning has been widely studied for its ability to learn policies through static datasets. However, the limited radiation range of static datasets leads to overestimation of some out-of-distribution (OOD) samples, causing suboptimal policy obtained through distribution shift - a critical challenge in offline RL.
While the policy regularization method addresses this by cloning dataset sampling policies, its effectiveness is restricted by excessive constraints on policy improvement.
To address this and further improve the performance of algorithms, a hypercube policy regularization framework is proposed, which combines local exploration with policy regularization. This method alleviates the constraints of policy regularization methods by allowing the agent to explore the actions corresponding to similar states in the static dataset. Theoretical guarantees are established for performance enhancement. In addition, when integrated with TD3-BC and Diffusion-QL as TD3-BC-C and Diffusion-QL-C, state-of-the-art results are achieved on D4RL benchmarks: a 24% improvement is observed in suboptimal datasets with significant performance gains across most environments.
@@
reinforcement learning, offline reinforcement learning, local exploration, policy regularization.
@@ 1 Introduction
The objective of Reinforcement Learning (RL) is to learn an optimal policy through the interaction between the agent and the environment. At present, there have been some notable achievements in the field, such as games[19], Large Model[16] and navigator[9].
@@
However, a potential issue with general RL is that the agent must interact with the environment during training, potentially causing harm to the environment or the agent, leading to economic losses[2]. One potential solution to this problem is offline Reinforcement Learning (offline RL)[13]. In offline RL, the agent tries to learn an optimal policy from a static dataset. Since the training of offline RL only requires a static dataset, it does not have the problem of the environment or agent being damaged during training. Nevertheless, the static dataset is typically unable to cover the full spectrum of potential state-action combinations, and the uncovered pairs are termed out-of-distribution (OOD) state-action pairs[6]. Q or V-value estimators may overestimate the value of some suboptimal OOD pairs, leading to a preference for selecting such OOD samples with high value and resulting in suboptimal policies. This challenge, known as the distribution shift problem, is a critical problem for offline RL.
@@
To address this issue, Current methods can be mainly classified into policy regularization[17], Q-value regularization[7], importance sampling[15], uncertainty estimation[3] and model-based methods[14]. Among the aforementioned methods, policy regularization methods attempt to clone the policy used for static datasets. Current algorithms utilizing policy regularization methods achieve good results when combined with advanced policy clone models such as diffusion models to clone the behavior of policy distribution[20]. However, the excessive constraints imposed by these methods severely limit the policy improvement, especially in low-quality datasets. When applied to low-quality static datasets, policies trained through this approach demonstrate inferior performance compared with other methodologies.
Q-value regularization algorithms achieve better policies by exploring OOD state-action pairs. However, unrestricted exploration requires precise Q-value constraints. The associated computational overhead significantly reduces practical applicability.
@@
The core limitation stems from an unresolved trade-off: Policy regularization ensures stability but sacrifices exploration capability, while Q-value regularization enables exploration but incurs prohibitive computational costs. Motivated by the complementary characteristics of the above two methods, the combination of policy regularization and Q-value regularization methods may yield superior results. Therefore, this article seeks to permit the agent to explore some promising actions and replicate the overall state actions that have been explored. It can be noted that for any given state, there may be multiple similar states within the static dataset. Furthermore, in datasets, especially low-quality datasets, such similar states may have better actions due to the randomness of the sampling actions. Therefore, an attempt is made to set the exploration scope of the agent in a given state limited to actions that correspond to adjacent states.
@@
Inspired by Latin Hypercube Sampling’s[18] efficient space partitioning, in the hypercube policy regularization framework, hypercubes are defined to allow for localized action exploration within state neighborhoods supported by the offline dataset. Compared with current policy regularization methods, which are ineffective in low-quality datasets, and the current Q-value regularization methods, which require a long training time[12]. The hypercube policy regularization framework directly addresses the limitations above, improves the performance of algorithms while maintaining the same training time as policy regularization methods. Another advantage of this framework is its versatility, which allows it to be applied to a variety of policy regularization algorithms such as TD3-BC[5] and Diffusion-QL[20]. The combined algorithm, TD3-BC-C and Diffusion-QL-C exhibited remarkable performance in experiments conducted on the D4RL dataset[4]. TD3-BC-C and Diffusion-QL-C can achieve better results compared to those of state-of-the-art(SOTA) algorithms such as Diffusion-QL.
@@
The remainder of this paper is organized as follows. Firstly, the related work is introduced. Secondly, a detailed explanation of the hypercube policy regularization framework is provided. Finally, experiments for the new algorithms are conducted on the D4RL dataset.
@@ 2 Related Work
This article is mainly related to the policy regularization method in offline RL. The basic idea of this method is to solve the distribution shift problem by constraining the distance between the training policy and the corresponding state-action pairs within the static dataset.
@@
Among the policy regularization methods, BCQ[6] considers the extrapolation error and adds disturbances based on the use of VAE to generate actions to solve the extrapolation error. BEAR[11] considers that previous algorithms cannot fit OOD state-actions well. To deal with it, BEAR adds weighted behavior clone loss in the optimization process of the policy to solve the problem. Compared with BCQ and BEAR, which require more operations, TD3-BC[5] achieves better results by adding only a few modifications to BC. Compared with the above algorithms, IQL [10] additionally introduces the state value function to update the Q-value function and policy. Diffusion-QL[20] achieves a state-of-the-art (SOTA) level by using conditional diffusion models for behavior clones and policy regularization. However, as mentioned before, policy regularization methods rely on data within static datasets, which may result in suboptimal results when the quality of static datasets is poor.
@@
On the other hand, advanced Q-value constraint algorithms constrain the Q-values of OOD state-actions and thus allow the agent to explore the entire state-action space. For instance, EDAC[1] and PBRL [3] can have good results in low-quality datasets. However, such algorithms usually require a long computing time to find an excellent policy.
Considering the characteristics of the above two methods, local space exploration is integrated into the overall policy regularizations. The exploration capacity of policy regularization algorithms is enhanced through the proposed approach, thereby improving algorithmic performance while reducing dependency on static dataset quality.
@@ 3 Proposed Method
In this section, the hypercube policy constraint framework is introduced, and a theoretical analysis of its performance is conducted. Subsequently, two novel algorithms are formulated through the integration of the proposed framework with baseline offline reinforcement learning (RL) algorithms.
@@ 3.1 Hypercube policy regularization framework
The hypercube constraint framework proposed in this section partitions the state space into hypercubes, where the exploration of any state-action pair is confined to similar states and corresponding actions within the same hypercube. The hypercube partitioning can be obtained through the following two steps. The initial step is to divide the state space in the static dataset into a hypercube by utilizing an integer $\delta$. This is achieved by obtaining the coordinates of the $i$-th dimension of any state $s$ in the static dataset in hypercube space, as outlined below:
@@
Figure 1: Hypercubed state space of dimension 3, choosing $\delta=2$.
Figure 2: All states can explore $a_{max}$ in the hypercube.
A hypercubed state space with dimension 3, the $\delta=2$ is illustrated in Fig 2. Each circle represents a distinct state within the static dataset. The number of states within each hypercube $\geq 0$. During the training phase, the agent is permitted to explore the actions corresponding to the states in the same cube.
@@
In detail, in any state, in order to select the better action and achieve the goal of enabling the agent to explore within a certain range, the Q-value $Q(s,a_{\max})$ of selecting the current optimal action within the same hypercube is compared with the Q-value $Q(s,a)$ of selecting the action corresponding to that state itself like Fig2. Fig 2 shows the hypercube v1 in Fig 2, in which any state action is allowed to explore the current optimal action $a_{max}$, thereby achieving local exploration. Finally, the hypercube policy regularization framework can be derived by integrating the constraint with a Markov decision process. Fig 3 illustrates a training flow chart that excludes the evaluation phase. At the beginning of training, the static dataset is hypercubed and used to constrain the current action $a$ to obtain a new action $a_{\rm new}$ during training. At the end of a loop, the process is reversed to update the policy $\pi_{\theta}$, and this step is repeated to output a policy $\pi_{\theta}$.
@@
3.2 Theoretical analysis
The hypercube policy constraint framework is guaranteed to maintain or enhance algorithm performance when Q-function approximators provide perfect value estimations. However, practical implementations must address imperfect Q-value estimations. In this section, theoretical analysis demonstrates that modifying the value of $\delta$ allows the hypercube constraint framework to enhance or sustain the performance of the algorithm regardless of the Q-value estimation.
Assumption 1
@@ 3.2 Theoretical analysis
$\left|{Q({s_{1}},{a_{1}})-Q({s_{2}},{a_{2}})}\right|\leq K_{Q}\left|{({s_{1}}\oplus{a_{1}})-({s_{2}}\oplus{a_{2}})}\right|$ (4)
Where $\oplus$ is the vector concatenation operation, this assumption can be satisfied when the Q-function is approximated by neural networks[8]. Thus, it could be demonstrated that the hypercube policy regularization framework ensures the algorithm’s performance does not deteriorate under Assumption 1.
Theorem 3.1.
@@
$Q(s,{a^{\pi_{\rm new}}})\geq Q(s,{a^{\pi_{\rm old}}})$ (5)
Proof 3.2.
When a single state-action pair $(s,a)$ exists in the hypercube, then $\pi_{\rm new}=\pi_{\rm old}$ and $Q(s,a^{\pi_{\rm old}})=Q(s,a^{\pi_{\rm new}})$.
When there is another state-action $(s^{\prime},a^{\prime})$ in the hypercube, $Q(s,a^{\prime})$ is discussed through $Q(s,a)$ and $Q(s^{\prime},a^{\prime})$. For this, three cases must be considered: $Q(s^{\prime},a^{\prime}), Q(s,a)=Q(s,a)$$Q(s^{\prime},a^{\prime})$ and $Q(s,a)>Q(s^{\prime},a^{\prime})$.
When $Q(s^{\prime},a^{\prime}), Q(s,a)Q(s,a)$ can be expressed as:
@@
$\begin{array}[]{c}Q(s,a^{\prime})\leq Q(s^{\prime},a^{\prime})+K|(s\oplus a^{\prime})-(s^{\prime}\oplus a^{\prime})|\\ =Q(s^{\prime},a^{\prime})+K|s-s^{\prime}|\end{array}$ (7)
Combining (6) and (7), it can be known that when $K|s-s^{\prime}|<b_{1}$ , $Q(s,a^{\prime})\leq Q(s,a)$.
@@
This guarantees that when $\theta>c_{1}$, $Q(s,a)>Q(s^{\prime},a^{\prime})$, the new policy is always worse than the original policy, so $\pi_{\rm new}=\pi_{\rm old}$, $Q(s,a^{\pi_{\rm new}})=Q(s,a^{\pi_{\rm old}})$ in this situation.
When $Q(s,a)=Q(s^{\prime},a^{\prime})$, $\pi_{\rm new}=\pi_{\rm old}$ is selected to obtain a conservative policy. Therefore, in this case, $Q(s,a^{\pi_{\rm new}})=Q(s,a^{\pi_{\rm old}})$.
@@
$\begin{array}[]{c}Q(s,a^{\prime})\geq Q(s^{\prime},a^{\prime})-K_{Q}|(s\oplus a^{\prime})-(s^{\prime}\oplus a^{\prime})|\\ =Q(s^{\prime},a^{\prime})-K_{Q}|s-s^{\prime}|\end{array}$ (10)
Combining (9) and (10), it can be known that when $K_{Q}|s-s^{\prime}|<b_{2}$, $Q(s,a^{\prime})\geq Q(s,a)$. Similar to $Q(s^{\prime},a^{\prime})<Q(s,a)$, there exists an integer $c_{2}$, when $\theta>c_{2}$,
$K_{Q}|s-s^{\prime}|\leq K_{Q}S_{max}\leq b_{2}$ (11)
This demonstrates that when $\theta>c_{2}$, $Q(s^{\prime},a^{\prime})>Q(s,a)$. Consequently, by designating $a^{\prime}$ as $a_{\rm new}$, the new policy can guarantee that $Q(s,a^{\pi_{\rm new}})\geq Q(s,a^{\pi_{\rm old}})$.
For the situation that there is only one extra $(s^{\prime},a^{\prime})$ in the hypercube and $\theta>max\{c_{1},c_{2}\}$. When $Q(s^{\prime},a^{\prime})\leq Q(s,a)$, the policy remains unchanged, $Q(s,a^{\pi_{\rm new}})=Q(s,a^{\pi_{\rm old}})$. When $Q(s^{\prime},a^{\prime})>Q(s,a^{\prime})$, $Q(s,a^{\pi_{\rm new}})\geq Q(s,a^{\pi_{\rm old}})$. Therefore, (5) holds in this case.
When there are multiple state-actions $(s_{i},a_{i}),i=1,2,...,n$ within the static dataset in each hypercube, refer to the above analysis of the existence of only one additional state-action in the hypercube, it means that When the Q-value of some state-actions is smaller than $Q(s,a)$, there exist $b_{1}>0$, $\theta>c_{1}$ make:
@@
Therefore, when there are multiple state-actions in a hypercube, selecting $\theta>max\{c_{1},c_{2}\}$ can also make $Q(s,{a^{\pi_{\rm new}}})\geq Q(s,{a^{\pi_{\rm old}}})$.
According to Therom3.1, a large one $\delta$ can ensure that the performance of the algorithms remains unaltered or improves. In practice, the Q-function approximator can produce a relatively accurate estimate of the state-action pairs within the static dataset. Consequently, the uncertainty associated with the Q-value of actions within the static dataset is relatively low. Therefore, a smaller $\delta$ should be used to make the agent explore more actions and thus achieve the best result.